OpenStudy (anonymous):

If a football player passes a football from 4 feet off the ground with an initial velocity of 36 feet per second, how long will it take the football to hit the ground? Use the equation h = –16t2 + 36t + 4. Using the Quadratic Formula answer please

OpenStudy (anonymous):

–16t^2 + 36t + 4 = 0 -4 (4t^2 - 9t - 1) = 0 (4t^2 - 9t - 1) = 0

OpenStudy (anonymous):

0 = -16t^ + 36t + 4 16t^2 - 36t - 4 = 0 4t^2 - 9t - 1 = 0 solve for t - ignore negative root

OpenStudy (anonymous):

this wont factorise so you'll need to use quadratic formula